OpenStudy (anonymous):

HELP ME ANSWER IT! explain! Which of the following is equivalent to a slope of negative four? a) slope equals negative four over negative one b) slope equals negative one over negative four c) slope equals negative four over positive one d) slope equals positive one over negative four

OpenStudy (darkprince14):

slope = rise over run

OpenStudy (anonymous):

I know that

OpenStudy (darkprince14):

a positive slope goes from lower left to upper right and vice versa.... a negative slope goes from upper left to lower right and vice versa...

OpenStudy (darkprince14):

that goes for the slope of a line:)

OpenStudy (anonymous):

ok so..

OpenStudy (darkprince14):

You will only get a negative slope if and only if either the numerator (rise) OR the denominator (run) is negative... Only one of them is negative...

OpenStudy (anonymous):

its b?

OpenStudy (darkprince14):

no, only one of them is negative, the other one is positive..

OpenStudy (anonymous):

so it has to be the one that says negative and its not b so its d?

OpenStudy (darkprince14):

If you look at it \[\frac{ -y }{ x } = -\left( \frac{ y }{ x } \right) = \frac{ y }{ -x }\] Is this equality right?

OpenStudy (anonymous):

yes it is.

OpenStudy (darkprince14):

so the answer could either be C or D:)

OpenStudy (anonymous):

its d right

OpenStudy (darkprince14):

the graph of a line could either go from left to right or right to left. There's no distinct direction. So the two of them can be correct. Solve C and Solve D. You'll get Negative four on both of them:)

OpenStudy (anonymous):

I got d

OpenStudy (darkprince14):

if that's your choice:)

OpenStudy (anonymous):

yes but is that correct because if I get this wrong I fail

OpenStudy (darkprince14):

Both of them are right so you'll pass don't worry. Sorry if i'm not that much of a help...

OpenStudy (anonymous):

you are good thanks